# An “opinion reproduction number” for infodemics in a bounded-confidence content-spreading process on networks

### v1

@@ III A content-spreading process with a bounded-confidence mechanism
We study a content-spreading model with a bounded-confidence mechanism for adopting an opinion and sharing content. However, in our model, receptive agents simply adopt the opinion that is espoused by content, rather than compromising their opinion as in a standard BCM.
@@
The bounded-confidence models in section II describes the time-dependent opinions of the agents of a network. However, in real life, one does not usually have easy access to numerical values of peoples opinions. One instead typically observes the content that is spread on social networks. Naturally, such content often reflects an underlying opinion or ideological belief. There is much empirical work on quantifying the dissemination and virality of content on social-media platforms like $\mathbb{X}$ (which is the platform formerly known as Twitter) Kümpel, Karnowski, and Keyling (2015); Muhammed T and Mathew (2022), and branching processes provide a natural model of content spread on social media. See, for example, Lerman and Ghosh (2010); Guille et al. (2013); Gleeson et al. (2014); Vosoughi, Roy, and Aral (2018); Gleeson et al. (2021).
@@
We seek to track the spread of fixed content with a fixed ideology (i.e., opinion) $x_{0}$. In other words, the spreading process has no mutation of ideas or editorializing. In the context of social media, one can interpret the fixed-content assumption as accounts only reposting or sharing content, without any additions or changes. By contrast, real social-media systems have a rich ecosystem of content evolution Adamic et al. (2016).
@@
As described in section II, we model a social network as a simple, undirected graph $G(V,\mathcal{E})$ with $|V|=N$ nodes. Let $A$ denote the graph’s associated $N\times N$ adjacency matrix. The adjacency-matrix entry $A_{ij}=1$ if there is an edge between nodes $i$ and $j$; otherwise, $A_{ij}=1$. The spread of content with opinion $x_{0}$ begins at a source node, which we select uniformly at random from $V$ and label as node 00. We initialize all other nodes $i\in\{1,\ldots,N-1\}$ in $V$ with a state $x_{i}$ that we draw from a distribution $\phi(x)$. We require $\phi(x)$ to be Riemann integrable.
@@
Given an opinion space $\Omega$, we refer to the opinion $x_{0}\in\Omega$ of the content as the content state. Additionally, we refer to the state $x_{i}\in\Omega$ of node $i\in\{1,\ldots,N-1\}$ as its opinion. In the present paper, we focus on $\Omega=[0,1]$, although in Figure 10 we show a situation with $\Omega=\mathbb{R}$. Let $c\in[0,1]$ denote the receptiveness parameter, which is akin to the confidence bound of a typical BCM. In practice, in the present paper, we consider $c\in(0,1/2]$, so that agents are not receptive to agents with overly different opinions. The spread of content with opinion $x_{0}$ begins at node 00. The content spreads through a percolation-inspired process Sahimi (2023) via an update rule that is based on a bounded-confidence mechanism (see algorithm 1). In fig. 1, we show a schematic of our algorithm on a small network.
@@
Input: model parameters $c,\ x_{0},\ \phi(x),\ N$, and the distribution of the graph degree sequence
@@
$x_{j}=\begin{cases}x_{i}\text{ (and $i$ is appended to {next nodes})}\,,&\text{if }|x_{j}-x_{0}|<c\\ x_{j}\,,&\text{otherwise}\end{cases}$
@@
Figure 1: A schematic illustration of our content-spreading process (see Algorithm 1). In the top row, we show consecutive update steps of Algorithm 1. We initialize a graph at time $t=0$ with one active node (in pink), which we label as node 00. We suppose that the root node 00 has content state $x_{0}=0.1$. We initialize all other nodes $j$ to have opinions $x_{j}\in(0,1)$. In this example, the receptiveness parameter is $c=0.35$. At each time step, any neighbors $j$ of the previously active node that satisfy $|x_{j}-x_{0}|<c$ become active (i.e., turn pink). We thus add them to the dissemination tree $G_{0}$. In this example, Algorithm 1 terminates after 4 steps. In the bottom row, we show the resulting dissemination tree $G_{0}$ in orange, with the root node outlined.
To study content spread on a network, we construct a dissemination tree $G_{0}$ with root node 00 to track the spreading process Goel et al. (2016); Oh and Porter (2018). Content spreads as follows. Suppose that content state $x_{0}$ starts at node 00. We look at the set $\mathcal{N}_{i}$ of neighbors $j$ of node $i$ and select all $j\in\mathcal{N}_{i}$ that satisfy $|x_{j}-x_{0}|<c$. These neighbors of $i$ spread the content (i.e., they become “activated” with the content and change their opinion $x_{j}$ to be equal to the content state $x_{0}$) if the distance between their opinion $x_{j}$ and the content state $x_{0}$ is sufficiently small. We represent this activation by adding these nodes to the dissemination tree $G_{0}$ as child nodes of node $i$. We then continue this process for each child node of the current dissemination tree until there are no available neighbors that are left to activate. In other words, we continue this process until either all nodes are activated or until there are no remaining neighbors of an active node.
Using Algorithm 1, it is guaranteed that $G_{0}$ is a tree because a node activates the first time that it encounters the content of one of its neighbors and can only be activated once. When two activated nodes share a receptive neighbor, we choose the activated node with the smaller index as the parent node for the newly activated neighbor. Modifications of this algorithm can result in a directed acyclic graph instead of a tree. It is worthwhile to examine such modifications in future work (especially if considering the effects of competing social contagions Min and San Miguel (2018)).
@@ IV Forecasting an “infodemic” on a large network
We now examine content spread on a network with infinitely many nodes. This situation approximates content spread on networks with many nodes (i.e., for large $N$), particularly in the early stages of content spread.
We follow the analysis in Noël et al. (2009), who examined a model of infectious-disease spread. We use generating functions to write down the expected number of neighbors that spread the source node’s content. We refer to this expected number of neighbors as the opinion reproduction number $R$. The opinion reproduction number, whose critical value $R=1$ is its associated infodemic threshold, is akin to the basic reproduction number $R_{0}$ in models of disease spread Kiss, Miller, and Simon (2017); Brauer, Castillo-Chavez, and Feng (2019).
@@ IV.1 Analysis
Suppose that content spreads on a configuration-model network Fosdick et al. (2018) with degree distribution $p_{k}$. We generate each configuration-model network using stub matching, without removing self-edges or multi-edges. Let $q_{k}$ denote its associated excess degree distribution. The generating function $g_{0}$ of the degree distribution and the generating function $g_{1}$ of the excess degree distribution $q_{k}$ are
@@
Given a node with $k$ neighbors, we want to write down the probability that $l$ of these $k$ neighbors spread a piece of content. To be able to do this, we need to calculate the probability that content is spread along one edge. This single-edge transmission probability depends on both the content state $x_{0}$, the receptiveness parameter $c$, and the distribution of the initial opinions $x_{i}$ (with $i\in\{1,\ldots,N\}$).
Suppose that we draw the initial opinions $x_{i}$ from a distribution with probability density function $\phi(x)$. The probability of content spread (i.e., transmission) along an edge depends on the probability of drawing a value $x_{i}$ that lies within distance $c$ of $x_{0}$. The single-edge transmission probability is
@@
One common choice of $\phi(x)$ is to draw the initial node opinions from the uniform distribution on the interval $(0,1)$. The probability of drawing an opinion $x_{i}$ within distance $c$ of $x_{0}$ is then $2c$, unless $x_{0}$ is within $c$ of the boundaries of $(0,1)$. This yields
@@
Another common choice of $\phi(x)$ is to draw the initial node opinions from a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$ (where the opinion space is $\mathbb{R}$). This yields
@@
Using the single-edge transmission probability [see eq. 3], we see that the probability that $l$ of $k$ neighbors spread the content is
@@
To write down the probability generating function for the number of second neighbors (i.e., the neighbors of neighbors) of a source node in a dissemination tree, we use a similar argument, but now we use the excess degree distribution instead of the degree distribution. We obtain
@@
If a network $G$ is infinite (or at least large enough so that we can neglect finite-size effects), nodes cannot be activated more than once, and the content spread yields a directed acyclic graph (i.e., the dissemination tree), it follows that $g_{a}\left(1+(z-1)s(x_{0},c)\right)=g_{1}\left(1+(z-1)s(x_{0},c)\right)$ is the probability generating function for the number of nodes that spread a piece of content that they received directly from a single node at level $a$ of the dissemination tree. Importantly, we are relying on the assumption that the number of times that a piece of content is spread is much smaller than the total number of nodes of a network. We expect our approximation to be reasonable in the first few generations of a dissemination tree, which describe the early stages of content spread.
In the onset of content dissemination, the expected number of “grandchildren” (i.e., second neighbors) in a dissemination tree that starts from a node that spreads the content at level $a$ is the opinion reproduction number
@@
where the prime ${}^{\prime}$ denotes differentiation with respect to $z$. When the number of content shares increases from one level to the next, the opinion reproduction number $R$ (which is analogous to the basic reproduction number $R_{0}$ in models of infectious-disease spread Kiss, Miller, and Simon (2017)) is larger than $1$. Similarly, a decrease in number of content shares from one level to the next corresponds to $R<1$. The critical value $R=1$ is the infodemic threshold. This calculation is reminiscent of branching-processes approaches that examine whether or not content spread tends to locally magnify or contract with time Gleeson et al. (2021).
@@
$R=s(x_{0},c)\frac{\lambda^{2}}{\lambda}=s(x_{0},c)\lambda$
because $g_{0}(z)=e^{-\lambda(1-z)}$. The transition value that determines whether content spread increases locally or decreases locally is
$R=\begin{cases}\frac{1}{\lambda}+x_{0}\,,&x_{0}\leq c\\ \frac{1}{2\lambda}\,,&c<x_{0}<1-c\\ \frac{1}{\lambda}-1+x_{0}\,,&x_{0}\geq 1-c\,.\end{cases}$
When the receptiveness parameter $c>R$, the expected number of content shares increases from one level to the next in the early stages of content spread, so we expect the content to take hold in the population. Unsurprisingly, the condition $c>R$ is satisfied most easily for larger receptive parameter $c$ and larger mean degree $\lambda$. The content state $x_{0}$ has less impact than $c$ and the node degree, except when the content state is near 00 or $1$.
@@ IV.2 Simulations
We now compare our analytical results from section IV.1 to numerical simulations of our content-spread model. We simulate the model on 10000-node configuration-model networks with degree sequences that we determine using a Poisson distribution with mean $\lambda=5$. We draw the initial opinions uniformly at random from the interval $(0,1)$. In fig. 2, we show a scenario in which our analysis indicates that we are above the infodemic threshold (i.e, $R>1$). In fig. 3, we show a scenario in which our analysis indicates that we are below the infodemic threshold (i.e., $R<1$). In this computation, the mean total number of content shares is very small. (It is about 10.)
@@
In fig. 4, we show a phase diagram that summarizes when our content-spreading model experiences experiences an infodemic for different values of the content state $x_{0}$ and receptiveness parameter $c$. We show the infodemic threshold, which we obtain using the condition $R=1$, from our analysis in orange.
Figure 2: A comparison of numerical simulations of our content-spreading model with our analytical predictions in an infodemic situation (i.e., when $R>1$). For each numerical simulation, we generate a 10000-node configuration-model network with a degree sequence that we determine using a Poisson distribution with mean $\lambda=5$. We draw the node opinions $x_{i}$ uniformly at random from $(0,1)$, and we set the content state to be $x_{0}=0.5$ and the receptiveness parameter to be $c=0.2$. In each realization, we choose $5$ source nodes uniformly at random to seed with content state $x_{0}$. The solid purple curve gives the total number of content shares up to and including time $t$ averaged over $100$ realizations. In each realization, we draw a new degree sequence and a new set of node opinions. The light purple shaded region indicates the standard deviation for these $100$ realizations. The dashed orange curve gives our analytical approximation.
@@
Figure 3: A comparison of numerical simulations of our content-spreading model with our analytical predictions in a situation without an infodemic (i.e., when $R<1$). For each numerical simulation, we generate a 10000-node configuration-model network with a degree sequence that we determine using a Poisson distribution with mean $\lambda=5$. We draw the node opinions $x_{i}$ uniformly at random from $(0,1)$, and we set the content state to be $x_{0}=0.5$ and the receptiveness parameter to be $c=0.05$. In each realization, we choose $5$ source nodes uniformly at random to seed with content state $x_{0}$. The solid purple curve gives the total number of content shares up to and including time $t$ averaged over $100$ realizations. In each realization, we draw a new degree sequence and a new set of node opinions. The light purple shaded region indicates the standard deviation for these $100$ realizations. The dashed orange curve gives our analytical approximation.
@@
Figure 4: The $(x_{0},c)$ phase diagram for 10000-node configuration-model networks with degree sequences that we generate using a Poisson distribution with mean $\lambda=5$. We vary both the content state ($x_{0}$, horizontal axis) and the receptiveness parameter ($c$, vertical axis). The shading of each square indicates the mean, across 100 trials, of the total number of content shares as a proportion of the total number of nodes, with lighter shades indicating that more nodes share the content. In each realization, we choose $5$ source nodes uniformly at random to seed with content state $x_{0}$. In each realization, we draw a new degree sequence and a new set of node opinions. The orange solid curve is the predicted opinion reproduction threshold $R$ from our analysis.

### v2

@@ III A content-spreading process with a bounded-confidence mechanism
We study a content-spreading model with a bounded-confidence mechanism for adopting an opinion and sharing content. However, in our model, receptive agents simply adopt the opinion that is espoused by content, rather than compromising their opinions as in a standard BCM.
@@
The bounded-confidence models that we discussed in Section II describe the time-dependent opinions of the agents of a network. However, in real life, one usually does not have easy access to numerical values of people’s opinions. One instead typically observes the content that is spread on social networks. Naturally, such content often reflects an underlying opinion or ideological belief. There is much empirical work on quantifying the dissemination and virality of content on social-media platforms like $\mathbb{X}$ (i.e., the platform formerly known as Twitter) Kümpel, Karnowski, and Keyling (2015); Muhammed T and Mathew (2022), and branching processes provide a natural model of content spread on social media. See, for example, Refs. [Lerman and Ghosh, 2010; Guille et al., 2013; Gleeson et al., 2014; Vosoughi, Roy, and Aral, 2018; Gleeson et al., 2021].
@@
We seek to track the spread of fixed content with a fixed opinion $x_{0}$. In other words, the spreading process has no mutation of ideas or editorializing. In the context of social media, one can interpret the fixed-content assumption as accounts only sharing or reposting content, without any additions or changes. By contrast, real social-media systems have a rich ecosystem of content evolution Adamic et al. (2016).
@@
As we described in Section II, we model a social network as an undirected, unweighted, and simple graph $G(V,\mathcal{E})$ with $|V|=N$ nodes. The spread of content with opinion $x_{0}$ begins at a source node, which we select uniformly at random from $V$ and label as node 00. We initialize all other nodes $i\in\{1,\ldots,N-1\}$ of a graph $G$ with a state $x_{i}$ that we draw from a distribution $\phi(x)$. In our analytical calculations, we assume either that there is a single source node or that the number of source nodes is sufficiently smaller than the network size $N$. We can then treat spreading processes that start from distinct nodes as independent of each other. We require $\phi(x)$ to be Riemann integrable.
@@
Given an opinion space $\Omega$, we refer to the opinion $x_{0}\in\Omega$ of the content as the content state. We refer to the state $x_{i}\in\Omega$ of node $i\in\{1,\ldots,N-1\}$ as its opinion. In the present paper, we focus on $\Omega=[0,1]$, although in Figure 9 we show a situation with $\Omega=\mathbb{R}$. Let $c\in[0,1]$ denote the receptiveness parameter, which is akin to the confidence bound of a typical BCM. In practice, in the present paper, we consider $c\in(0,1/2]$, so agents are not receptive to agents with overly different opinions. The spread of content with opinion $x_{0}$ begins at node 00. The content spreads through a percolation-inspired process Sahimi (2023) via an update rule that is based on a bounded-confidence mechanism (see Algorithm 1). In Figure 1, we show a schematic of our algorithm on a small network.
@@
Input: Model parameters $c$, $x_{0}$, $\phi(x)$, $N$ and the degree distribution
@@
$x_{j}=\begin{cases}x_{i}\text{ {(add $i$ to {next nodes})}}\,,&\text{if }|x_{j}-x_{0}|<c\\ x_{j}\,,&\text{otherwise}\end{cases}$
@@
Figure 1: A schematic illustration of our content-spreading process (see Algorithm 1). In panel (a), we show several consecutive update steps of Algorithm 1. We initialize a graph at time $t=0$ with one active node (in pink), which we label as node 00. We suppose that the root node 00 has content state $x_{0}=0.1$. We initialize all other nodes $j$ to have opinions $x_{j}\in(0,1)$. In this example, the receptiveness parameter is $c=0.35$. At each time step, any neighbors $j$ of the previously active node that satisfy $|x_{j}-x_{0}|<c$ become active (i.e., turn pink). We thus add them to the dissemination tree $G_{0}$. In this example, Algorithm 1 terminates after 4 steps. In panel (b), we show the resulting dissemination tree $G_{0}$ in orange, with the root node outlined in black.
To study content spread on a network, we construct a dissemination tree $G_{0}$ with root node 00 to track the spreading process Goel et al. (2016); Oh and Porter (2018). Content spreads as follows. Suppose that content state $x_{0}$ starts at node 00. We look at the set $\mathcal{N}_{i}$ of neighbors $j$ of node $i$ and select all nodes $j\in\mathcal{N}_{i}$ with an opinion $x_{j}$ that satisfies $|x_{j}-x_{0}|<c$. These neighbors of $i$ spread the content (i.e., they “activate” and change their opinions $x_{j}$ to the content state $x_{0}$) because the distance between the opinion $x_{j}$ and the content state $x_{0}$ is sufficiently small. We represent this activation by adding these nodes to the dissemination tree $G_{0}$ as child nodes of node $i$. We then continue this process for each child node of the current dissemination tree until there are no available neighbors that are left to activate. In other words, we continue this process until either all nodes are activated or until no active nodes have any remaining receptive neighbors.
Using Algorithm 1, it is guaranteed that $G_{0}$ is a tree because a node activates the first time that it encounters the content of one of its neighbors and can only activate once. When two active nodes share a receptive neighbor, we choose the active node with the smaller index as the parent node of the newly activated neighbor. Modifications of this algorithm can result in a directed acyclic graph instead of a tree. It is worthwhile to examine such modifications in future work (e.g., if considering the effects of competing social contagions Min and San Miguel (2018)).
@@ IV Forecasting an “infodemic” on a large network
We now examine content spread on a network with infinitely many nodes. This situation approximates content spread on networks with many nodes (i.e., large $N$), particularly in the early stages of content spread.
We use analysis that was developed for models of infectious-disease spread on networks Newman, Strogatz, and Watts (2001); Newman (2002); Noël et al. (2009). We use generating functions to obtain an expression for the expected number of neighbors that spread the source node’s content. We refer to this expected number of neighbors as the opinion reproduction number $R$. The opinion reproduction number, whose critical value $R=1$ is its associated infodemic threshold, is akin to the basic reproduction number $R_{0}$ in models of disease spread Kiss, Miller, and Simon (2017); Brauer, Castillo-Chavez, and Feng (2019).
@@ IV.1 Analysis
Suppose that content spreads on a configuration-model network Fosdick et al. (2018) with degree distribution $p_{k}$. We generate each configuration-model network by uniformly randomly matching “stubs” (i.e., ends of edges) to each other. We remove all self-edges and multi-edges after matching stubs, so different networks can have slightly different degree sequences. Let $q_{k}$ denote the associated excess degree distribution. The generating function $g_{0}$ of the degree distribution and the generating function $g_{1}$ of the excess degree distribution $q_{k}$ are
@@
Given a node with $k$ neighbors, we want to determine the probability that $l$ of these $k$ neighbors spread a piece of content. To do this, we need to calculate the probability that content is spread along one edge. This single-edge transmission probability depends on the content state $x_{0}$, the receptiveness parameter $c$, and the distribution of the initial opinions $x_{i}$ (with $i\in\{1,\ldots,N\}$).
Suppose that we draw the initial opinions $x_{i}$ from a distribution with probability density function $\phi(x)$. The probability of content spread (i.e., transmission) along an edge depends on the probability of drawing a value $x_{i}$ that lies within a distance $c$ of $x_{0}$. The single-edge transmission probability is
@@
One common choice of $\phi(x)$ is to draw the initial node opinions from the uniform distribution on the interval $(0,1)$. The probability of drawing an opinion $x_{i}$ within a distance $c$ of $x_{0}$ is then $2c$, unless $x_{0}$ is within $c$ of the boundaries of $(0,1)$. This yields
@@
Another common choice of $\phi(x)$ is to draw the initial node opinions from a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$. (In this case, the opinion space is $\mathbb{R}$.) This yields
@@
Using the single-edge transmission probability [see Equation 3], we see that the probability that $l$ of $k$ neighbors spread the content is
@@
To obtain the probability generating function for the number of second neighbors (i.e., the neighbors of neighbors) of a source node in a dissemination tree, we use a similar argument, but now we use the excess degree distribution instead of the degree distribution. This yields the probability generating function
@@
If a network $G$ is infinite (or at least large enough so that we can neglect finite-size effects), nodes cannot activate more than once, and the content spread yields a directed acyclic graph (i.e., a dissemination tree). Consequently, $g_{a}\left(1+(z-1)s(x_{0},c)\right)=g_{1}\left(1+(z-1)s(x_{0},c)\right)$ is the probability generating function for the number of nodes that spread a piece of content that they received directly from a single node at level $a$ of a dissemination tree. Importantly, we are relying on the assumption that the number of times that a piece of content is spread is much smaller than the total number of nodes of a network.
In the early stages of content dissemination, the expected number of “grandchildren” (i.e., second neighbors) in a dissemination tree that starts from a node that spreads the content at level $a$ is the opinion reproduction number
@@
where the prime denotes differentiation with respect to $z$. The right-hand side of Equation 9 relates the expected number of grandchildren to the expected number of children. When the number of content shares increases from one level to the next, the opinion reproduction number $R$ (which is analogous to the basic reproduction number $R_{0}$ in models of infectious-disease spread Kiss, Miller, and Simon (2017)) is larger than $1$. Similarly, a decrease in the number of content shares from one level to the next corresponds to $R<1$. The critical value $R=1$ is the infodemic threshold. This calculation is reminiscent of branching-processes approaches that examine whether or not content spread tends to locally magnify or contract with time Gleeson et al. (2021). Given an opinion reproduction number $R$, we approximate the expected number of content shares up to and including time step $T$ as $n_{0}\sum_{t=0}^{T}R^{t}$, where $n_{0}$ is the number of source nodes and we treat content that spreads from different source nodes as independent. Given our assumptions, we expect our approximation to be reasonable in the first few generations of a dissemination tree. These initial generations describe the early stages of content spread.
@@
$R=s(x_{0},c)\frac{\lambda^{2}}{\lambda}=s(x_{0},c)\lambda$ (10)
because $g_{0}(z)=e^{-\lambda(1-z)}$. Substituting Equation 4 into Equation 10 allows us to calculate the opinion reproduction number in terms of the content state $x_{0}$, the receptiveness parameter $c$, and $\lambda$. At the infodemic threshold $R=1$, the critical value $c^{*}$ of the receptiveness parameter that determines whether content spread increases locally or decreases locally is
$c^{*}=\begin{cases}\frac{1}{\lambda}+x_{0}\,,&x_{0}\leq c^{*}\\ \frac{1}{2\lambda}\,,&c^{*}<x_{0}\leq 1-c^{*}\\ \frac{1}{\lambda}-1+x_{0}\,,&x_{0}>1-c^{*}\,.\end{cases}$
When the receptiveness parameter $c>c^{*}$, the expected number of content shares increases from one level to the next in the early stages of content spread, so we expect the content to take hold in a network. Unsurprisingly, the infodemic threshold $R=1$ is exceeded most easily for larger values of the receptiveness parameter $c$ and networks with larger expected mean degree $\lambda$. The content state $x_{0}$ has less impact than the receptiveness parameter $c$ and the expected mean degree, except when the content state is near 00 or $1$.
@@ IV.2 Simulations
We now compare our analytical results from Section IV.1 to numerical simulations of our content-spreading model. We simulate the model on 10000-node configuration-model networks with degree sequences that we determine using a Poisson distribution with mean $\lambda=5$. We draw the initial opinions uniformly at random from the interval $(0,1)$. In Figure 2(a), we show a scenario in which our analysis indicates that we are above the infodemic threshold (i.e, $R>1$). In Figure 2(b), we show a scenario in which our analysis indicates that we are below the infodemic threshold (i.e., $R<1$). In this computation, the mean total number of content shares is very small. (It is about $10$.)
@@
In Figure 3, we show a phase diagram that summarizes when our content-spreading model experiences an infodemic for different values of the content state $x_{0}$ and receptiveness parameter $c$. In orange, we show the critical value $c^{*}$ that we obtain analytically using the infodemic threshold $R=1$.
@@
Figure 2: A comparison of numerical simulations of our content-spreading model with our analytical predictions in (a) an infodemic situation (when the receptiveness parameter is $c=0.2$, which implies that $R>1$) and (b) a situation without an infodemic (when the receptiveness parameter is $c=0.05$, which implies that $R<1$). For each numerical simulation, we generate a 10000-node configuration-model network with a degree sequence that we determine using a Poisson distribution with mean $\lambda=5$. We draw the initial node opinions $x_{i}$ uniformly at random from $(0,1)$, and we set the content state to $x_{0}=0.5$. In each realization, we choose $5$ source nodes uniformly at random to seed with content state $x_{0}$. In our analytical calculations, we assume that the content spread from each source node is independent, as the number of source nodes is much smaller than the total number of nodes. The solid purple curve gives the total number of content shares (averaged over $100$ realizations) as a function of time. In each realization, we draw a new degree sequence and a new set of node opinions. The light purple shaded region indicates the standard deviation for these $100$ realizations. The dashed orange curve gives our analytical approximation.
@@
Figure 3: The $(x_{0},c)$ phase diagram for 10000-node configuration-model networks with degree sequences that we generate using a Poisson distribution with mean $\lambda=5$. We vary both the content state $x_{0}$ and the receptiveness parameter $c$. The shading of each square indicates the mean, across 100 trials, of the total number of content shares as a proportion of the total number of nodes, with lighter shades indicating that more nodes share the content. In each realization, we choose $5$ source nodes uniformly at random to seed with content state $x_{0}$. In each realization, we draw a new degree sequence and a new set of node opinions. The orange solid curve shows the critical receptiveness value $c^{*}$ for each $x_{0}$.